# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v1

@@ II.4 Green’s functions, momentum distributions, and densities
$G_{\mathrm{F}}^{-1}\left(\textbf{k},\omega\right)$ $=G_{\mathrm{F}}^{0}\left(\textbf{k},\omega\right)^{-1}-\Sigma_{F}\left(\textbf{k},\omega\right),$ (12)
@@
The three equations (16-18) are at the core of our study. They constitute a system of non-linear integral equations whereby the condensate density and the bosonic and fermionic chemical potentials are solved for, once the values of the densities $n_{\mathrm{B}},n_{\mathrm{F}}$ and scattering lengths $a_{\mathrm{BF}}$ and $a_{\mathrm{BB}}$ are given.
III Asymptotic limit of strong boson-fermion coupling
In this section we will examine the regime of strong boson-fermion coupling ${\varepsilon_{0}}/{E_{\mathrm{F}}}\equiv 2\text{e}^{2g}\gg 1$ where a simplified and more transparent approach can be adopted. The system of Eqs. (16-18) is replaced with an equivalent but much simpler one, for which a semi-analytical solution can be found. This allows us to disclose the microscopic physics underlying the different concentration regimes ($0\leq x\leq 1$) in the strong coupling regime ($g>2$), and, at the same time, to perform stringent checks on crucial physical quantities otherwise obtained through a full numerical implementation. The direct boson-boson repulsion is neglected in the present regime, since it is expected to produce negligible effects.
III.1 Boson-Fermion propagator $T$(${\mathbf{P}}$, $\Omega$)
We first delve into the nature of the molecular states through an analysis of the pair propagator $T({\mathbf{P}},\Omega)$. In the limit ${\varepsilon_{0}}/{E_{\mathrm{F}}}\gg 1$, the two contributions in Eq. (5) to the $T$-matrix $\Gamma({\mathbf{P}},\Omega)$ acquire a simplified form. The first one (see Appendix A, Eq. (66)) reduces to the polar form
@@ III.1 Boson-Fermion propagator $T$(${\mathbf{P}}$, $\Omega$)
with $\mu_{\mathrm{CF}}\equiv\mu_{\mathrm{F}}+\mu_{\mathrm{B}}+{\varepsilon_{0}}$ the chemical potential of the composite fermions, whereas the second contribution (see Appendix A, Eq. (67)) stems from the atomic fermions and reduces to
$I_{\mathrm{F}}({\mathbf{P}},\Omega)\approx\int\frac{d\textbf{p}}{\left(2\pi\right)^{2}}\frac{\Theta\left(-\xi_{\textbf{P}-\textbf{p}}^{F}\right)}{{\varepsilon_{0}}}=\frac{n_{\mu_{\mathrm{F}}}^{0}}{{\varepsilon_{0}}},$ (20)
where $n_{\mu_{\mathrm{F}}}^{0}$ is the number density of a non-interacting Fermi gas with chemical potential ${\mu_{\mathrm{F}}}$. When Eqs. (19) and (20) are inserted into Eq. (5), the latter acquires the strong coupling form
@@
which suggests the presence of an effective self-energy term $\Sigma_{\mathrm{CF}}=2\pi n_{\mu_{\mathrm{F}}}^{0}/m_{r}$. The latter represents an effective repulsion field generated by the medium of atomic fermions (${\mu_{\mathrm{F}}}>0$) onto the existing molecules, and is an effective way to take into account the kinetic energy cost of populating the molecular Fermi sphere in addition to the atomic one. Through the constant form factor $2\pi{\varepsilon_{0}}/m_{r}$, stemming from the composite nature of the molecules, the above $T$-matrix $\Gamma_{\mathrm{SC}}({\mathbf{P}},\Omega)$ ends up being proportional to a bare fermionic Green’s function $G_{\rm CF}^{0}({\mathbf{P}},\Omega)^{-1}=i\Omega-P^{2}/2M+{\mu_{\mathrm{CF}}}-\Sigma_{\mathrm{CF}}$ dressed by the mean-field of the atomic medium.
In 2D the three-body $T$-matrix in vacuum, when treated within the Born approximation, reduces to a constant value $2\pi/m_{r}$, implying that $\Sigma_{\mathrm{CF}}$ is a Hartree contribution of the interaction between the molecular and atomic species (see Appendix B). The atom-molecule scattering length is actually expected to vanish in the strong-coupling limit of the BF interaction [97]. The coupling-independent atom-molecule scattering length obtained by the present approach thus overestimates the atom-molecule repulsion.
Being a mere mean-field shift, $\Sigma_{\mathrm{CF}}$ can be absorbed into a redefinition of the chemical potential in Eq. (19), ${\mu_{\mathrm{CF}}}\rightarrow{\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}-\Sigma_{\mathrm{CF}}$, thus leading to
@@
Atomic fermions can also correlate with bosons belonging to the condensate during a pair-breaking event (see Fig. 1(b)); this process is taken into account in the dressed propagator $T({\mathbf{P}},\Omega)$ as a self-energy insertion in the related Dyson’s (or Bethe-Salpeter) equation (7) which in the strong coupling limit acquires the form
$T_{\mathrm{SC}}^{-1}\left(\textbf{P},\Omega\right)=\Gamma_{\mathrm{SC}}^{-1}\left(\textbf{P},\Omega\right)-n_{0}G_{F}^{0}\left(\textbf{P},\Omega\right)$
@@
$\equiv\frac{m_{r}}{2\pi{\varepsilon_{0}}}G_{\mathrm{CF}}^{-1}({\mathbf{P}},\Omega)$ (24)
with $G_{\mathrm{CF}}({\mathbf{P}},\Omega)$ indicating the dressed composite fermion propagator.
The quantity $2\pi{\varepsilon_{0}}n_{0}/m_{r}\equiv\Delta_{0}^{2}$ naturally arises as an energy scale associated with exchange of bosons with the condensate and reveals its meaning through an analysis of the pole structure of $T_{\mathrm{SC}}({\mathbf{P}},\Omega)$, by which one finds that the resulting dispersion relations
$E_{\mathbf{P}}^{\pm}=\frac{\tilde{\xi}^{\mathrm{CF}}_{\mathbf{P}}+\xi^{\mathrm{F}}_{\mathbf{P}}\pm\sqrt{(\tilde{\xi}^{\mathrm{CF}}_{\mathbf{P}}-\xi^{\mathrm{F}}_{\mathbf{P}})^{2}+4\Delta_{0}^{2}}}{2}$ (25)
denote hybridization of a bare molecule (with dispersion ($\tilde{\xi}_{\mathbf{P}}^{\mathrm{CF}}=P^{2}/2M-{\tilde{\mu}_{\mathrm{CF}}}$) with an unpaired atom (with dispersion $\xi_{\mathbf{P}}^{\mathrm{F}}=P^{2}/2m_{\mathrm{F}}-{\mu_{\mathrm{F}}}$) and a boson in the condensate (with vanishing kinetic energy). This effect was originally pointed out in [62] for a two-channel model, and signals the inception of strongly renormalized molecular states. These states will be discussed at length in Sec. III.4 after the thermodynamic parameters will have been consistently determined.
The pair propagator $G_{\mathrm{CF}}({\mathbf{P}},\Omega)$ acquires a simple form when expressed in terms of the quasi-particle energies $E_{\mathbf{P}}^{\pm}$ and corresponding weights $u_{\mathbf{P}}^{2}$ and $v_{\mathbf{P}}^{2}$,
@@
As a consequence, the number of composite fermions $n_{\mathrm{CF}}$ is swiftly obtained for given coupling and population imbalance as
$n_{\mathrm{CF}}$ $=\int\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\frac{d\Omega}{2\pi}G_{\mathrm{CF}}({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}$
$=\int\frac{d{\mathbf{P}}}{(2\pi)^{2}}\left[u_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{+})+v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})\right].$ (28)
The above integral has a fully analytical form, which is given in Eq. (85) of appendix D. In practice, however, only the term containing $v_{\mathbf{P}}^{2}$ contributes, since $E_{\mathbf{P}}^{+}$ is always positive in the coupling range of interest here.
In conclusion, we have found that molecular BF states are: i) renormalized by the presence of the remaining unpaired fermions via the mean-field self-energy $\Sigma_{\mathrm{CF}}$, and ii) hybridized with atomic states owing to pair-breaking fluctuations that foster a non-vanishing condensate. For small bosonic concentrations, the latter feature can also be viewed as feedback of polaronic correlations on the two-body correlator $T({\mathbf{P}},\Omega)$.
III.2 Fermionic self-energy and momentum distribution
We now pass to examine the fermionic self-energy Eq. (11) in the strong coupling limit, that is with $T(\textbf{P},\Omega)$ replaced by $T_{\mathrm{SC}}(\textbf{P},\Omega)$
The frequency integration in Eq. (11) is computed via a contour integration on the left-hand side of the complex plane, owing to the presence of the causality factor $0^{+}$. Therefore, the pole originating from $G_{\mathrm{B}}^{0}({\mathbf{P}}-{\mathbf{k}},\Omega-\omega)$ does not contribute and one is left with the following expression
$\Sigma_{\mathrm{F}}({\mathbf{k}},\omega)$ $=\frac{\Delta^{2}_{0}}{i\omega-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}}+\left(\frac{2\pi{\varepsilon_{0}}}{m_{r}}\right)\int\frac{d{\mathbf{P}}}{(2\pi)^{2}}$
×$\times\left[{u_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{+})\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{P}}-{\mathbf{k}}}-E_{\mathbf{P}}^{+}}+{v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{P}}-{\mathbf{k}}}-E_{\mathbf{P}}^{-}}\right].$ (29)
Taking advantage of the fact that ${\mu_{\mathrm{B}}}\simeq-{\varepsilon_{0}}$ is the dominant energy scale in the denominators appearing in the integrand on the r.h.s. of Eq. (III.2), one is allowed to take ${\mathbf{P}}=0$ therein thus obtaining
@@ III.2 Fermionic self-energy and momentum distribution
where we have neglected $E_{{\mathbf{P}}=0}^{\pm}$ altogether with respect to ${\mu_{\mathrm{B}}}$ and where $n_{\mathrm{CF}}$ is given by Eq. (III.1).
@@
$G_{\mathrm{F}}^{-1}({\mathbf{k}},\omega)=i\omega-\xi_{\mathbf{k}}^{\rm F}-\frac{\Delta^{2}_{0}}{i\omega-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}}-{\Delta^{2}_{\mathrm{CF}}\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}},$ (32)
which has three poles in the complex plane. One pole is essentially given by $z=-\xi^{\rm B}_{{\mathbf{k}}}$ while the other two are roughly given by $z=\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}$ and $z=\xi_{\mathbf{k}}^{\rm F}$. An analogous form was obtained in [98, 99] even though in the normal phase and in 3D.
Before passing to the exact evaluation of Eq. (32) in closed form, it is instructive for its physical interpretation to analyze the limit of small $\Delta_{0}/{E_{\mathrm{F}}}$, which is expected to occur either when $x\to 1$ (depletion of the condensate density due to increase of molecule number) or $x\to 0$ (reduction of the condensate density due to decrease of boson number). This analysis will provide us with a clear decomposition of $n_{\mathrm{F}}({\mathbf{k}})$ into its atomic and molecular contributions. By expanding $G_{\mathrm{F}}({\mathbf{k}},\omega)$ to first order in $\Delta_{0}/{E_{\mathrm{F}}}$, one obtains
$G_{\mathrm{F}}({\mathbf{k}},\omega)$ $\approx{1\over i\omega-\xi_{\mathbf{k}}^{\rm F}-{\Delta^{2}_{\mathrm{CF}}\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}}}+$
$+\frac{\Delta_{0}^{2}}{\left(i\omega-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}\right)\left(i\omega-\xi^{\mathrm{F}}_{\mathbf{k}}-{2\pi n_{\mathrm{CF}}\over m_{r}}\right)^{2}}\ ,$ (33)
where we have further approximated
$\frac{\Delta_{\mathrm{CF}}^{2}}{i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}+\mu_{\mathrm{CF}}}\approx-\frac{\Delta_{\mathrm{CF}}^{2}}{\mu_{\mathrm{B}}}={2\pi n_{\mathrm{CF}}\over m_{r}}\equiv\Sigma_{\mathrm{F}}^{0}$ (34)
in the denominator of the second term. The quantity $2\pi n_{\mathrm{CF}}/m_{r}\equiv\Sigma_{\mathrm{F}}^{0}$ is analogous to $\Sigma_{\mathrm{CF}}$ of Sec. III.1 in that it represents the repulsion induced by the mean-field of the molecules onto the unpaired fermions and can be interpreted as a Hartree mean-field shift (${\tilde{\mu}_{\mathrm{F}}}\equiv{\mu_{\mathrm{F}}}-\Sigma_{\mathrm{F}}^{0}$).
Let us first consider the first term in Eq. (III.2). It can be recast in the more familiar (BCS-like) form
$\frac{\bar{u}^{2}_{k}}{i\omega-\bar{E}_{\mathbf{k}}^{+}}+\frac{\bar{v}^{2}_{k}}{i\omega-\bar{E}_{\mathbf{k}}^{-}}$ (35)
with
$\bar{E}_{\mathbf{k}}^{\pm}={1\over 2}\left(\xi_{\mathbf{k}}^{\mathrm{F}}-\xi^{\mathrm{B}}_{{\mathbf{k}}}\pm\sqrt{(\xi_{\mathbf{k}}^{\rm F}+\xi^{\mathrm{B}}_{{\mathbf{k}}})^{2}+4\Delta_{\mathrm{CF}}^{2}}\right)$ (36)
and
$\bar{u}^{2}_{k}$ $={1\over 2}\left(1+\frac{\xi_{\mathbf{k}}^{\rm F}+\xi^{\mathrm{B}}_{{\mathbf{k}}}}{\sqrt{(\xi_{\mathbf{k}}^{\rm F}+\xi^{\mathrm{B}}_{{\mathbf{k}}})^{2}+4\Delta_{\mathrm{CF}}^{2}}}\right)$ (37)
$\bar{v}^{2}_{k}$ $=1-\bar{u}^{2}_{k}.$ (38)
$\bar{E}_{\mathbf{k}}^{+}\simeq\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}},\quad\quad\quad\bar{E}_{\mathbf{k}}^{-}\simeq-\xi_{\mathbf{k}}^{\mathrm{B}},$ (39)
$\bar{u}^{2}_{k}\simeq 1-\frac{\Delta_{\mathrm{CF}}^{2}}{\left({{\mathbf{k}}^{2}\over 2m_{r}}+{\varepsilon_{0}}\right)^{2}},\quad\quad\bar{v}^{2}_{k}\simeq\frac{\Delta_{\mathrm{CF}}^{2}}{\left({{\mathbf{k}}^{2}\over 2m_{r}}+{\varepsilon_{0}}\right)^{2}}.$ (40)
having defined $\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}}\equiv\xi_{\mathbf{k}}^{\mathrm{F}}+\Sigma_{\mathrm{F}}^{0}$. Closing the integration contour on the left-hand side of the plane as prescribed by Eq. (14), from the first term in Eq. (III.2) one obtains the contribution
$n_{\mathrm{F}}^{(1)}({\mathbf{k}})=\left(1-\frac{{2\pi{\varepsilon_{0}}\over m_{r}}n_{\mathrm{CF}}}{\left({{\mathbf{k}}^{2}\over 2m_{r}}+{\varepsilon_{0}}\right)^{2}}\right)\Theta\left(-\tilde{\xi}_{\mathbf{k}}^{\rm F}\right)+\frac{{2\pi{\varepsilon_{0}}\over m_{r}}n_{\mathrm{CF}}}{\left({{\mathbf{k}}^{2}\over 2m_{r}}+{\varepsilon_{0}}\right)^{2}}$ (41)
where one immediately recognizes the squared amplitude of the two-body wave-function in vacuum

### v2

@@ 2.4 Green’s functions, momentum distributions, and densities
$G_{\mathrm{F}}^{-1}\left(\textbf{k},\omega\right)$ $=G_{\mathrm{F}}^{0}\left(\textbf{k},\omega\right)^{-1}-\Sigma_{\rm F}\left(\textbf{k},\omega\right),$ (13)
@@
The three equations (17)-(19) are at the core of our study. They constitute a system of non-linear integral equations whereby the condensate density and the bosonic and fermionic chemical potentials are solved for, once the values of the densities $n_{\mathrm{B}},n_{\mathrm{F}}$ and scattering lengths $a_{\mathrm{BF}}$ and $a_{\mathrm{BB}}$ are given.
3 Strong boson-fermion coupling regime
In this section, we will examine the regime of strong boson-fermion coupling ${\varepsilon_{0}}/{E_{\mathrm{F}}}\equiv 2\text{e}^{2g}\gg 1$, where a simplified and more transparent approach can be adopted. The system of Eqs. (17)-(19) is replaced with an equivalent but much simpler one, for which a semi-analytical solution can be found. This allows us to disclose the microscopic physics underlying the different concentration regimes ($0\leq x\leq 1$) in the strong-coupling regime ($g>2$), and, at the same time, to perform stringent checks on crucial physical quantities otherwise obtained through a full numerical implementation. The direct boson-boson repulsion is expected to produce only minor effects in this regime, which is dominated by BF pairing and the corresponding large binding energy. For this reason, we altogether neglect it in the present section.
3.1 Many-body $T$-matrix in the condensed phase $T$(${\mathbf{P}}$, $\Omega$)
We first delve into the nature of the molecular states by analyzing the many-body $T$-matrix in the condensed phase $T({\mathbf{P}},\Omega)$ which, as we will see, for strong boson-fermion coupling, acquires the meaning of a composite fermion propagator.
In the limit ${\varepsilon_{0}}/{E_{\mathrm{F}}}\gg 1$, the two contributions in Eq. (6) to $\Gamma({\mathbf{P}},\Omega)$ acquire a simplified form. The contribution $T_{2}({\mathbf{P}},\Omega)$ reduces to the polar form (see Appendix A, Eq. (A.11))
@@ 3.1 Many-body $T$-matrix in the condensed phase $T$(${\mathbf{P}}$, $\Omega$)
where $\mu_{\mathrm{CF}}\equiv\mu_{\mathrm{F}}+\mu_{\mathrm{B}}+{\varepsilon_{0}}$ and $M\equiv{m_{\mathrm{B}}}+{m_{\mathrm{F}}}$ correspond to the chemical potential and mass of the composite fermions, respectively.
The contribution $I_{\mathrm{F}}({\mathbf{P}},\Omega)$ can instead be approximated as (see Appendix A, Eq. (A.13))
$I_{\mathrm{F}}({\mathbf{P}},\Omega)\approx\frac{n_{\mu_{\mathrm{F}}}^{0}}{{\varepsilon_{0}}},$ (21)
where $n_{\mu_{\mathrm{F}}}^{0}={m_{\mathrm{F}}}{\mu_{\mathrm{F}}}\Theta(\mu_{\rm F})/2\pi$ is the number density of a non-interacting Fermi gas with chemical potential ${\mu_{\mathrm{F}}}$.
When Eqs. (20) and (21) are inserted into Eq. (6), the latter acquires the strong-coupling form
@@
which suggests the presence of an effective self-energy term $\Sigma^{0}_{\mathrm{CF}}=2\pi n_{\mu_{\mathrm{F}}}^{0}/m_{r}$. The latter represents an effective repulsion field generated by the medium of atomic fermions (${\mu_{\mathrm{F}}}>0$) onto the existing molecules and is an effective way to take into account the kinetic energy cost of populating the molecular Fermi sphere in addition to the atomic one.
In 2D the three-body $T$-matrix in vacuum, when treated within the Born approximation, reduces to a constant value $2\pi/m_{r}$, implying that $\Sigma^{0}_{\mathrm{CF}}$ is a Hartree contribution of the interaction between the molecular and atomic species (see Appendix B). The atom-molecule scattering length is actually expected to vanish in the strong-coupling limit of the BF interaction [108]. The coupling-independent atom-molecule scattering length obtained by the present approach thus overestimates the atom-molecule repulsion.
Being a mere mean-field shift, $\Sigma^{0}_{\mathrm{CF}}$ can be absorbed into a redefinition of the chemical potential in Eq. (20), ${\mu_{\mathrm{CF}}}\rightarrow{\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}-\Sigma^{0}_{\mathrm{CF}}$, thus leading to
@@
$\Gamma_{\mathrm{SC}}({\mathbf{P}},\Omega)$ then appears to be proportional to a bare-like fermionic Green’s function
${\tilde{G}}^{0}_{\mathrm{CF}}({\mathbf{P}},\Omega)\equiv\frac{1}{i\Omega-P^{2}/2M+{\tilde{\mu}_{\mathrm{CF}}}}$ (24)
of a particle of mass $M$ with bare dispersion $P^{2}/2M-\mu_{\rm CF}$ renormalized by a mean-field shift $\Sigma^{0}_{\mathrm{CF}}$ due to interaction with the atomic medium, the constant of proportionality $2\pi{\varepsilon_{0}}/m_{r}$ in Eq. (20) originating from the composite nature of the molecules.
Atomic fermions can also correlate with bosons belonging to the condensate during a pair-breaking event (see Fig. 1(b)); this process is taken into account in the many-body $T$-matrix in the condensed phase $T({\mathbf{P}},\Omega)$ as a self-energy insertion in the related Dyson’s (or Bethe-Salpeter) equation (8) which in the strong coupling limit acquires the form
$T_{\mathrm{SC}}^{-1}\left(\textbf{P},\Omega\right)$ $=\Gamma_{\mathrm{SC}}^{-1}\left(\textbf{P},\Omega\right)-n_{0}G_{F}^{0}\left(\textbf{P},\Omega\right)$ (25)
@@
$\equiv\frac{m_{r}}{2\pi{\varepsilon_{0}}}G_{\mathrm{CF}}^{-1}({\mathbf{P}},\Omega).$ (27)
Here, $G_{\mathrm{CF}}({\mathbf{P}},\Omega)$, as defined by Eq. (27), can be interpreted as a dressed composite fermion propagator, resulting from the hybridization between an undressed molecular bound state, described by ${\tilde{G}}^{0}_{\mathrm{CF}}({\mathbf{P}},\Omega)$, and an unpaired state made of a fermionic atom and a boson belonging to the condensate.
Specifically, the quantity
$\Delta_{0}^{2}\equiv 2\pi{\varepsilon_{0}}n_{0}/m_{r}$ (28)
naturally arises as an energy scale associated with the exchange of bosons with the condensate and reveals its meaning through an analysis of the pole structure of $T_{\mathrm{SC}}({\mathbf{P}},\Omega)$, by which one finds the dispersion relations
$E_{\mathbf{P}}^{\pm}=\frac{\tilde{\xi}^{\mathrm{CF}}_{\mathbf{P}}+\xi^{\mathrm{F}}_{\mathbf{P}}\pm\sqrt{(\tilde{\xi}^{\mathrm{CF}}_{\mathbf{P}}-\xi^{\mathrm{F}}_{\mathbf{P}})^{2}+4\Delta_{0}^{2}}}{2}.$ (29)
They denote the hybridization of a molecule with renormalized dispersion $\tilde{\xi}_{\mathbf{P}}^{\mathrm{CF}}=P^{2}/2M-{\tilde{\mu}_{\mathrm{CF}}}$ with an unpaired atom with dispersion $\xi_{\mathbf{P}}^{\mathrm{F}}=P^{2}/2m_{\mathrm{F}}-{\mu_{\mathrm{F}}}$ and a boson in the condensate with vanishing kinetic energy. This kind of hybridization was first found in [65] within a two-channel model.
One sees from Eq. (29) that $\Delta_{0}$ controls the amount of hybridization between $\tilde{\xi}_{\mathbf{P}}^{\mathrm{CF}}$ and $\xi_{\mathbf{P}}^{\mathrm{F}}$ and that, for small $\Delta_{0}$, the hybridized dispersions $E_{\mathbf{P}}^{\pm}$ tend either to $\tilde{\xi}_{\mathbf{P}}^{\mathrm{CF}}$ or $\xi_{\mathbf{P}}^{\mathrm{F}}$ depending on the sign of $\tilde{\xi}^{\mathrm{CF}}_{\mathbf{P}}-\xi^{\mathrm{F}}_{\mathbf{P}}$, as it will be detailed in section 3.4. We note in this respect that the scale $\Delta_{0}$ vanishes when the product ${\varepsilon_{0}}n_{0}\to 0$. If, like in 3D, $n_{0}$ would vanish identically above a certain coupling strength for all $x\leq 1$, then one would obtain $\Delta_{0}=0$ in the strong-coupling limit for all concentrations considered in the present work. However, as already mentioned, $n_{0}$ vanishes only exponentially with the coupling strength in our approach, and the presence of the exponentially large factor ${\varepsilon_{0}}$ in the definition of $\Delta_{0}$ can compensate for the vanishing of $n_{0}$. We will see indeed below that, in the strong-coupling limit, $\Delta_{0}\to 0$ only when $x\to 0$ or $x\to 1$, with $\Delta_{0}$ tending to a finite (non-vanishing) value for all intermediate concentrations.
The composite-fermion propagator $G_{\mathrm{CF}}({\mathbf{P}},\Omega)$ acquires a simple form when expressed in terms of the quasi-particle energies $E_{\mathbf{P}}^{\pm}$ and corresponding weights $u_{\mathbf{P}}^{2}$ and $v_{\mathbf{P}}^{2}$,
@@
The number of composite fermions $n_{\mathrm{CF}}$ is swiftly obtained for given coupling and population imbalance as
$n_{\mathrm{CF}}$ $=\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\!\!\!\frac{d\Omega}{2\pi}G_{\mathrm{CF}}({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}$ (32)
$=\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\left[u_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{+})+v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})\right]$ (33)
where the integrand in Eq. (33) can be interpreted as the composite-fermion momentum distribution function $n_{\mathrm{CF}}\left({\mathbf{P}}\right)$. We further note that, in practice, only the term containing $v_{\mathbf{P}}^{2}$ contributes in Eq. (33) since $E_{\mathbf{P}}^{+}$ turns out to be always positive for the concentrations and couplings considered in our work. One thus has
$n_{\mathrm{CF}}\left({\mathbf{P}}\right)\equiv\left[u_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{+})+v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})\right]=v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-}).$ (34)
We finally note that the integration over ${\mathbf{P}}$ in Eq. (33) can be performed in a closed form (see Eq. (D.1) of appendix D).
In conclusion, we have found that molecular BF states are: i) renormalized by the presence of the remaining unpaired fermions via the mean-field self-energy $\Sigma^{0}_{\mathrm{CF}}$, and ii) hybridized with atomic states owing to pair-breaking fluctuations that foster a non-vanishing condensate. For small bosonic concentrations, the latter feature can also be viewed as feedback of polaronic correlations on the many-body $T$-matrix $T({\mathbf{P}},\Omega)$.
3.2 Fermionic self-energy and momentum distribution
We now pass to examine the fermionic self-energy Eq. (12) in the strong-coupling limit, that is, with $T(\textbf{P},\Omega)$ replaced by $T_{\mathrm{SC}}(\textbf{P},\Omega)$.
The frequency integration in Eq. (12) is computed via a contour integration on the left-hand side of the complex plane, owing to the presence of the convergence factor $e^{i\omega 0^{+}}$. The pole originating from $G_{\mathrm{B}}^{0}({\mathbf{P}}-{\mathbf{k}},\Omega-\omega)$ is located on the right-hand side of the complex plane because $\mu_{\rm B}<0$, and does not contribute to the integral. The contributions from the poles of $T_{\mathrm{SC}}$ yield instead
$\Sigma_{\mathrm{F}}({\mathbf{k}},\omega)=\frac{\Delta^{2}_{0}}{i\omega-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}}+\left(\frac{2\pi{\varepsilon_{0}}}{m_{r}}\right)\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\left[{u_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{+})\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{P}}-{\mathbf{k}}}-E_{\mathbf{P}}^{+}}+{v_{\mathbf{P}}^{2}\Theta(-E_{\mathbf{P}}^{-})\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{P}}-{\mathbf{k}}}-E_{\mathbf{P}}^{-}}\right].$ (35)
In the strong-coupling limit and for $x\leq 1$, the boson chemical potential ${\mu_{\mathrm{B}}}\simeq-{\varepsilon_{0}}$ is the dominant energy scale. One is then allowed to take ${\mathbf{P}}=0$ in the denominators of the integrand of Eq. (35) and neglect $E_{{\mathbf{P}}=0}^{\pm}$, which is of order $E_{\rm F}$, with respect to ${\mu_{\mathrm{B}}}$ therein, thus obtaining
@@ 3.2 Fermionic self-energy and momentum distribution
where $n_{\mathrm{CF}}$ is given by Eq. (34).
@@
$G_{\mathrm{F}}^{-1}({\mathbf{k}},\omega)=i\omega-\xi_{\mathbf{k}}^{\rm F}-\frac{\Delta^{2}_{0}}{i\omega-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}}-{\Delta^{2}_{\mathrm{CF}}\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}}.$ (38)
An analogous form was obtained in [109, 110], albeit in the normal phase and in 3D.
The Green’s function (38), when integrated over the frequency as in Eq. (15), determines the fermion momentum distribution $n_{\rm F}({\mathbf{k}})$. A simple contour integration then yields
where $z_{i},\;i=1,2,3$ are the three different (real) roots of the cubic equation
$z-\xi_{\mathbf{k}}^{\mathrm{F}}-\frac{\Delta^{2}_{0}}{z-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}}-{\Delta^{2}_{\mathrm{CF}}\over z+\xi^{\mathrm{B}}_{{\mathbf{k}}}}=0.$ (40)
One root is essentially given by $z=-\xi^{\rm B}_{{\mathbf{k}}}$, while the other two are roughly given by $z=\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}$ and $z=\xi_{\mathbf{k}}^{\rm F}$.
A more explicit expression for $n_{\mathrm{F}}({\mathbf{k}})$ can be obtained when the hybridization energy $\Delta_{0}\ll{E_{\mathrm{F}}}$. Within our theoretical approach, this condition occurs either when $x\to 1$ or when $x\to 0$. One obtains (details are provided in Appendix D.2)
$n_{\mathrm{F}}({\mathbf{k}})={n_{\mathrm{CF}}}|\phi({\mathbf{k}})|^{2}+n_{\mathrm{UF}}({\mathbf{k}}),$ (41)
where